OpenStudy (anonymous):

Convert this number hexadecimal into binary number : 1. 2A5C 2.C09 3.45FE

OpenStudy (anonymous):

MUSTARD

OpenStudy (anonymous):

2 A 5 C 0010 1010 0101 1 1 00

OpenStudy (anonymous):

good job SENSEI

OpenStudy (amistre64):

hex is by 16s right?

OpenStudy (anonymous):

yes right

OpenStudy (amistre64):

2 10 5 12 16^3 16^2 16^1 16^0 --------------------- 8192 + 2560 +80+12 = 10844 in decimal

OpenStudy (anonymous):

C 0 9 1100 0000 1001

OpenStudy (anonymous):

4 5 F E 0100 0101 1111 1110

OpenStudy (anonymous):

Easy way to convert from Hex to binary take one bit of Hex, and write as 4 bit binary

OpenStudy (anonymous):

sorry amistre hexa to binary

OpenStudy (anonymous):

look at table hexa to binary

OpenStudy (amistre64):

10844/2 = 5422 R 0 5422/2 = 2711 R 0 2711/2 = 1355 R 1 1355/2 = 677 R 1 677/2 = 338 R 1 i was working on it :)

OpenStudy (amistre64):

mighta gotten lost

OpenStudy (amistre64):

11100 so far 338/2 = 169 R 0 169/2 = 84 R 1 84/2 = 42 R 0 42/2 = 21 R 0 21/2 = 10 R 1 10/2 = 5 R 0 5/2 = 2 R 1 2/2 = 1 R 0 1010 0101 1100 maybe?

OpenStudy (anonymous):

There is a snappier way to do it

OpenStudy (amistre64):

im missing a 10 on the front ...

OpenStudy (amistre64):

oh i forgot to take it one more time

OpenStudy (amistre64):

1/2 = 0 R 1 lol

OpenStudy (amistre64):

10 1010 0101 1100 :)

OpenStudy (anonymous):

imran right